OpenStudy (anonymous):

Find the values of the six trigonometric functions of theta Function value Constraint sin theta= 3/5 theta lies in quadrant 2 PLEASE TELL ME HOW TEST TOMORROW

OpenStudy (bibby):



OpenStudy (bibby):

since sin = opposite/hypotenuse you know that: the opposite side is 3 the hypotenuse is 5. using the Pythagorean theorem you find the bottom side is 4

OpenStudy (bibby):

now using SOH-CAH-TOA you can find the rest to be cos= 4/5 tan=3/4 csc= 5/3 sec= 5/4 cot = 4/3